### v1

@@ 10 Proofs
In this section, we present the full proof of Theorem 1 to certify the global optimally regarding the solved local optima. Additionally, we present the proof of Remark 1 as a special case with bounded hyper-rectangle search space only given the Lipschitz constant of the objective function.
10.1 Full Proof of Theorem 1
Theorem 10.1 (Optimality Certification)
@@ 10.1 Full Proof of Theorem 1
which concludes the proof. $\blacksquare$
10.2 Proof of Remark 1
Remark 2
When $\mathcal{U}$ is a hyper-rectangle with the bounded $\ell_{2}$ norm of domain $U_{i}\in\mathbb{R}$ at each dimension $i$, with $i=1,2,\dots,n$, we remark Theorem 10.1 can hold in a more general way by only assuming that the Lipschitz constant $k_{G}$ of the objective function is given, where the following optimality certification regarding $x\in\mathcal{U}$ holds that:
@@ 10.2 Proof of Remark 1
Proof
Similar to the proof of Thm 10.1, by the sampling over subset $S\subset\mathcal{U}$ with $\ell_{2}$-norm density $\delta$, we have:
@@
which concludes the proof. $\blacksquare$
11 Discussion
@@
11.1 Limitations
11.1.1 Data Collection.
@@ 11.1 Limitations
11.1.2 LiDAR Configurations.
@@
11.1.3 Surrogate Metric.
@@
11.2 Potential Societal Impact
LiDAR systems, by their very nature, are designed to capture detailed information about the environment. This can include not just the shape and location of objects, but potentially also capturing point cloud data of individuals, vehicles, and private property in high resolution. The data captured might be misused if not properly regulated and secured. Further, With autonomous vehicles navigating spaces using LiDAR and other sensors, there could be a shift in how spaces are designed, potentially prioritizing efficiency for autonomous vehicles over human-centered design principles. In addition, over-reliance on LiDAR and perception systems can lead to questions about the trustworthiness and reliability of these systems,
12 Public Resources Used
In this section, we acknowledge the use of the following public resources, during the course of this work:

### v2

@@ Appendix E E Proofs
In this section, we present the full proof of Theorem 1 to certify the global optimally regarding the solved local optima. Additionally, we present the proof of Corollary 1 as a special case with bounded hyper-rectangle search space only given the Lipschitz constant of the objective function.
E.1 Full Proof of Theorem 1
Theorem 2 (Optimality Certification).
@@ E.1 Full Proof of Theorem 1
which concludes the proof. ∎
E.2 Proof of Corollary 1
Corollary 2.
When $\mathcal{U}$ is a hyper-rectangle with the bounded $\ell_{2}$ norm of domain $U_{i}\in\mathbb{R}$ at each dimension $i$, with $i=1,2,\dots,n$, Thm.2 can hold in a more general way by only assuming that the Lipschitz constant $k_{G}$ of the objective function is given, where the following optimality certification regarding $x\in\mathcal{U}$ holds that:
@@ E.2 Proof of Corollary 1
Proof.
Similar to the proof of Thm. 2, by the sampling over subset $S\subset\mathcal{U}$ with $\ell_{2}$-norm density $\delta$, we have:
@@
which concludes the proof. ∎
Appendix F F Discussions
@@
F.1 Limitations
F.1.1 Possible Limitation in Data Collection
@@
F.1.2 Possible Limitation in LiDAR Configurations
@@
F.1.3 Possible Limitation in Surrogate Metric
@@
F.2 Potential Societal Impact
LiDAR systems, by their very nature, are designed to capture detailed information about the environment. This can include not just the shape and location of objects, but potentially also capturing point cloud data of individuals, vehicles, and private property in high resolution. The data captured might be misused if not properly regulated and secured. Further, With autonomous vehicles navigating spaces using LiDAR and other sensors, there could be a shift in how spaces are designed, potentially prioritizing efficiency for autonomous vehicles over human-centered design principles. In addition, over-reliance on LiDAR and perception systems can lead to questions about the trustworthiness and reliability of these systems.
Appendix G G Public Resources Used
We acknowledge the use of the following public resources, during the course of this work: